OpenStudy (anonymous):

If f(x)= 2/3x^2+8x, what is the value of f(6)?

OpenStudy (anonymous):

try to replace x with 6

OpenStudy (campbell_st):

substitute x = 6 and then calculate an answer it looks like \[\frac{2}{3 \times 6^2 + 8 \times 6}\] then simplify the fraction

OpenStudy (anonymous):

@campbell_st thanks.

OpenStudy (anonymous):

@campbell_st wouldn't the answer be 62?

OpenStudy (campbell_st):

no, it should be a fraction \[\frac{2}{108 + 48} \]

OpenStudy (campbell_st):

not just simplify this

OpenStudy (anonymous):

\[\frac{ 1 }{78}\]